THE C-N STRETCHING VIBRONIC BANDS OF THE MgNC Ã ^{2}Π - X̃ ^{2}Σ^{+} TRANSITION

\underline{MASARU FUKUSHIMA} and TAKASHI ISHIWATA, Faculty of Information Sciences, Hiroshima City University, Asa-Minami, Hiroshima 731-3194, Japan.

We have generated MgNC in supersonic free jet expansions, and measured the laser induced fluorescence excitation spectra of the C-N stretching vibronic bands of the Ã ^{2}Π - X̃ ^{2}Σ^{+} transition. Rotational analysis yields the molecular constants of the vibronic levels, (1,0,0) and (1,0,1), in the Ã ^{2}Π state. The vibronic levels, (0,0,0), (0,1,0), and (0,0,1), were already analyzed by Wright and Miller^{a}. We can not find any anomalies in the constants of the C-N stretching vibronic levels, while they are predicted to be lying above the barrier of the isomerization reaction pathway, MgNC ↔ MgCN, on the Ã 2Π state. On the basis of the molecular constants obtained, we discuss the fine structures of both the ground X̃ ^{2}Σ^{+} and excited Ã ^{2}Π states.

aR. R. Wright and T. A. Miller, J. Mol. Spectrosc. 194, 219 (1999).